Solve this problem.
Office space. The length of a professor`s office is x feet, and the width is x + 4 feet. Write a polynomial that represents the area. and find the area if x = 10ft.

Answer by wuwei96815(245)

In this situation area=length times width.

A=L times W
A=x times x+4
A=x2 + 4x

If x = 10"
A=(10)2 + 4(10) = 100 + 40 = 140 square feet

I hope that is right?